Wednesday 11/06/12

9:45 A – 8:30 P

  • write to ML about bistability
  • write to Bender about Hox model.  — apparently email contact is not recommended.  Try phone tomorrow.
  • write summary of Francis suggestions
  • remove medium from S2 cells, add fresh medium.  Reduce total volume to 50 mL.
  • write to Eric about MBNL
  • send in PO for latest fly order (PcG mutants, marked balancer 3, hs-en).
Meeting with Fred and Brian 
  • got 20kb Fab7 BAC probe with A488 from Fred (100 ng of probe.  Dilute in 30 uL of DFISH buffer with embryos)
  • Got probe template and labeled primer set from Brian and Nick.  50 kb region
  • Brian will send browser for designing probes (basically has pre-tiled Drosophila genome with ideal probes (repeat screened, TM optimized etc) as a BED file.  Simply intersect ROI BED file with this data set to get probe set.  


  • check sample.  Focus lossed O/N.  Reset focus.  should keep an eye on this, will probably continue to get lost.  Monitor with remote desktop.
  • Switching still looks similar to yesterday, Pyranose system working so far.
  • switching looking a little sparse, may be cell type.
  • lost focus again (<11:30A), system very sensitive 🙁 — maybe a double spot issue.
  • Getting substantially fewer localizations from 750 channel.  May be buffer issue, may be section issue?  Will let scope try out a few more sections before re-etching.
  • substantial number of localizatios in 647 channel not turning off properly.  really should expand linkage and throw out long-ON dots more.  Better filtering needed and should not be too hard.
  • 750 density becoming pretty crappy.  647 looks qualitatively good still but total localization may have dropped substantially.


This entry was posted in Summaries and tagged , , . Bookmark the permalink.